There are 5 white and 45 black marbles in an urn. Standing next to the urn,...

  1. There are 5 white and 45 black marbles in an urn. Standing next to the urn, you close your eyes and draw 10 marbles with replacement.

a. What is the probability that you draw exactly 4 white marbles?

b. What is the probability that you draw at least 6 white marbles?

2. Studies show that gasoline use for compact cars sold in the US is normally distributed with a mean of 25.5 miles per gallon (mpg) and a standard deviation of 4.5 mpg. What percentage of compacts get 30 mpg or more?
Answer #1

1) let X : be the event of drawing white marbles ( with replacement) which follows binomial distribution !

  Binomial distribution: P(X=x) = nCx px qn-x

P(white marble) = 5/(5+45) = 0.1 => p= 0.1 and q=1- p = 0.9 and n=10

a)  probability of exactly 4 white marbles = P(X=4)= 10C4 * (0.1)4 * (0.9)10-4 = 0.011160 [answer]

b) P(at least 6 white marbles) = P(X= 6) + P(X= 7) + P(X= 8) + P(X= 9) + P(X=10)

= 10C6 *(0.1)6 * (0.9)4 + 10C7 *(0.1)7 * (0.9)3 + 10C8 *(0.1)8 * (0.9)2 + 10C9 * (0.1)9 *(0.9) + 10C10 * (0.1)10 *(0.9)0

= 0.000147 [ answer]

2)  

Given that ,

mean = = 25.5

standard deviation = = 4.5

P(x 30) = 1 - P(x   30)

= 1 - P[(x - ) / (30 - 25.5) / 4.5]

= 1 -  P(z 1)  

= 1 - 0.8413

= 0.1587

Answer = 15.87%

Any doubts please ask !! and please rate !! thank you ! :))